Let is observed the height (cm) of the students in a school, and sample with size of 40 students is consider:
165,185,172,151,164,151,169,176,166,178,178,166,158,
165,151,172,184,174,158,172,164,151,169,176,166,178,
176, 158, 152,165,172,186,174,158,184,170,152,164,161,166.
After grouping the data into 6 class intervals with equal width:
a) Make a summary table for the grouped frequency, grouped relative frequency and “less than” grouped cumulative frequency distribution for the given data.
1
Expert's answer
2012-10-26T11:25:47-0400
Height (cm)& frequency relative frequency cumulative frequency
151 – 156& 6& & & & & & & & 6/40 = 3/20 & 6
157 – 162& 5& & 5/40 = 1/8& & & 6+5 = 11
163 – 168& 10& 10/40 = ¼ & 11 + 10 = 21
169 – 174& 9& & & & & & & & & & 9/40& & & & 21 + 9 = 30
175 – 180& 6& & 6/40 = 3/20& 30+6 = 36
181 – 186& 4& 4/40 = 1/10& 36 + 4 = 40